OpenStudy (anonymous):

need to get points to graph y-3x=0

OpenStudy (anonymous):

rearrange the linear equation as y = 3x now put x =0, so y= 3*0 ==> y=0 so first point is (0,0) now put x = 2, so y= 3*2 ==> y=6 so second point is (2,6) now put x = -2 so y= 3*-2 ==> y = -6 so third point is (-2, -6) plot these three points and join them with a line..... y get the required graph...... if u r happy with the solution, pls click GOOD ANSWER button for me.....

OpenStudy (anonymous):

i just need two points to graph. so ill take the first two. Thanks!

OpenStudy (anonymous):

i suggest you take the second and the third one.......

OpenStudy (anonymous):

ok